OpenStudy (anonymous):

What is the height of a triangle that has an area of 60 yd2 and a base with a length of 12 yd?

OpenStudy (ganpat):

Area of triangle = 1/2 * base * ht 60 = 1/2 * 12 * ht Ht = 10 yd...